# Flow Matching and Diffusion Models via PointNet for Generating Fluid Fields on Irregular Geometries

### v1

@@ 4.1 Overall results
In this section, we present a comprehensive analysis of the results using tables, histograms, and visual comparisons. Table 2 reports the relative pointwise errors ($L^{2}$ norm) for predicting the velocity components $u$ and $v$ and the pressure $p$ using the three models: Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. The results are computed across the test set, which contains 222 unseen geometries. Specifically, for each geometry in the test set, 100 samples are generated using both the Flow Matching PointNet and Diffusion PointNet models, meaning that 100 random noise realizations are conditioned on the same geometry. The reported values in Table 2 are expressed in the form $\alpha_{1}\pm\alpha_{2}$, where $\alpha_{1}$ denotes the mean error and $\alpha_{2}$ represents the corresponding standard deviation. According to the values listed in Table 2, the average relative pointwise errors ($L^{2}$ norm) over the 222 geometries for the $u$-velocity component are 1.40%, 1.26%, and 3.20% for the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models, respectively. For the $v$-velocity component, the corresponding errors are approximately 4.81%, 6.17%, and 13.28%. These results demonstrate that the Flow Matching and Diffusion models, particularly for the $v$ variable, achieve substantially lower prediction errors. In comparison with the baseline PointNet, they are capable of reducing the relative pointwise error to below 10%, which is a remarkable improvement from an engineering standpoint. Predicting the $v$ variable is more challenging than predicting $u$ because the flow field exhibits both positive and negative velocities in the $y$-direction due to vortex formation behind the cylinder. Although the $u$ variable may also become negative in some regions, its spatial gradients are typically less steep than those of the $v$ variable.
@@
Examining the results for the pressure variable ($p$) reveals a similar trend. Both the Flow Matching PointNet and Diffusion PointNet models reduce the average prediction error to below 10%, whereas the baseline PointNet exceeds this limit. According to Table 2, the relative pointwise errors ($L^{2}$ norm) for predicting the pressure field are approximately 3.67%, 3.29%, and 10.83% for the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, respectively. The accuracy of the pressure field is particularly important because it directly influences the computation of lift and drag forces, which will be discussed in detail in Sect. 4.5. Furthermore, the uncertainty quantification capabilities of the proposed generative Flow Matching PointNet and Diffusion PointNet models provide valuable insights for practical engineering applications where safety factors must be considered in design processes. Finally, it is worth emphasizing that, as reported in Table 2, the number of trainable parameters, and consequently the model size, is identical across all three frameworks.
@@
In addition to reporting the mean, minimum, and maximum errors across the 222 test geometries in Table 2, the distribution of these errors is illustrated in the histogram shown in Fig. 4. As observed, the Flow Matching PointNet and Diffusion PointNet models demonstrate notably higher success in predicting the $v$ and $p$ variables. However, in all three models, there exist a few outlier cases for which the prediction error increases noticeably, as can be seen in Fig. 4. When comparing the Flow Matching PointNet and Diffusion PointNet in terms of prediction accuracy, the difference between the two is not substantial. Based on the overall results presented in Table 2 and Fig. 4, the Flow Matching PointNet model predicts the $v$-velocity component on average about 1.3% more accurately than the Diffusion PointNet model.
@@
Fig. 5: Evolution of velocity and pressure fields during the reverse process of Flow Matching PointNet at $n=0$ (i.e., standard Gaussian noise), $n=500$, $n=800$, $n=950$, $n=980$, and $n=1000$ (i.e., the clean field).
@@ 4.2 Reverse generative process
Figures 5 and 6 respectively illustrate the reverse processes of the Flow Matching PointNet and the Diffusion PointNet for fluid flow fields around a cylinder with a rectangular cross section, shown as a representative example from the test set. More specifically, Figure 5 presents the velocity and pressure fields at different reverse steps, including $n=0$ (i.e., where all three fields are sampled from a standard Gaussian distribution), $n=500$, $n=800$, $n=950$, and $n=980$, culminating at $n=1000$, which corresponds to the predicted clean velocity and pressure fields. As can be seen from Fig. 5, at $n=500$, the overall structure of the velocity and pressure fields is still largely visually obscured; however, the magnitudes of the variables are approximately within the range of [3, 3][-3,\,3], compared to the case $n=0$, where the variables lie roughly in the range of [6, 6][-6,\,6] due to the standard Gaussian noise. Similarly, at $n=800$, the noise level decreases such that the flow patterns become partially visible. At $n=950$, the velocity and pressure fields are largely reconstructed, and during the remaining reverse steps, they are further refined, with the predicted values converging to physically acceptable ranges. Quantitative error metrics have already been reported in Table 2 and discussed in the previous paragraphs.
Figure 6 shows similar trends for the reverse process of the Diffusion PointNet, starting from $t=1000$ (i.e., the standard Gaussian noisy field) and progressing to $t=0$, which corresponds to the predicted velocity and pressure fields. By comparing the outputs shown in Figures 5 and 6, it can be observed that the underlying mechanisms of the Flow Matching PointNet and the Diffusion PointNet differ. For example, consider the velocity and pressure field predictions obtained by the Flow Matching PointNet at $n=950$ and by the Diffusion PointNet at $t=50$, corresponding to the same number of reverse steps in the two models. Given the results shown in Fig. 5 and Fig. 6, a visual comparison indicates that the fluid flow patterns obtained by the Flow Matching PointNet exhibit a lower noise level and more closely resemble the predicted clean flow fields at this step (i.e., $n=950$) than those obtained by the Diffusion PointNet at the corresponding step (i.e., $t=50$). This qualitative observation highlights a fundamental difference in the reverse dynamics of the two modeling approaches.
@@ 4.3 Visual evaluation
Following the numerical analysis, a visual comparison of the results is conducted. Figures 710 present four examples of the Flow Matching PointNet model predictions for velocity and pressure fields, alongside the corresponding ground truth and absolute pointwise error distributions. These figures include cylinders with pentagonal, rectangular, and elliptical cross-sections. Similarly, Figures 1114 illustrate the results of the Diffusion PointNet model for cylinders with hexagonal, triangular, rectangular, and elliptical cross-sections. In all cases, the predictions demonstrate excellent accuracy and a strong agreement with the ground truth. The majority of the absolute errors occur on or near the cylinder surface, where the geometric variations among different test cases are most significant. In Fig.9 and Fig. 13, the $x$ component of the velocity vector ($u$) becomes negative, indicating the formation of secondary flow regions, and both the Flow Matching PointNet and Diffusion PointNet models successfully capture these phenomena. It should be noted that the test set is identical for all three models: Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. For visual diversity, however, we intentionally select cylinders with different cross-sectional shapes to provide a more comprehensive visual representation in Figs. 714. These results highlight the models’ success in handling diverse geometries with varying shapes, orientations, and sizes. Another notable observation is that the computational domain size differs for each geometry. For example, as can be seen in Fig. 7, the approximate minimum and maximum $x$-positions are 5 m and 11 m, while the corresponding $y$-positions range from 13 m to 19 m. In contrast, as shown in Fig. 9, the $x$-positions vary approximately from 2 m to 14 m, and the $y$-positions from 10 m to 22 m. A similar comparison can be made between Fig. 11 and Fig. 14. This adaptability reflects a fundamental strength of the PointNet architecture, which can process irregular geometries with varying domain sizes, unlike conventional CNNs that require fixed domain dimensions across the entire dataset. Additionally, the prediction of the noise derivative (by Flow Matching PointNet) or the noise (by Diffusion PointNet) at each point depends on both the spatial coordinates of that point and the global geometry of the point cloud that contains it. Since the spatial ranges of the $x$ and $y$ coordinates vary across different point clouds in the dataset, accurately predicting the noise derivative or the noise is a challenging task. The results presented here demonstrate, for the first time, that the proposed Flow Matching PointNet and Diffusion PointNet models successfully handle this challenge.
@@ 4.4 Sampling diversity
As discussed in Sect. 3, when random noise conditioned on a given geometry (point cloud) is provided as input to the Flow Matching PointNet and Diffusion PointNet models, they generate distinct realizations of the velocity and pressure fields. Here, we compare several of these generated samples. Figure 15 presents five samples generated by the Flow Matching PointNet model. To better illustrate the differences among the generated samples, instead of visualizing the velocity and pressure fields directly, we show the absolute pointwise error along with the corresponding ground truth. As can be seen from Fig. 15, it is evident that each sample differs from the others, even though the regions with the highest errors appear in roughly the same locations. However, the maximum absolute error for the $y$ component of the velocity vector ($v$) and the pressure field ($p$) still varies noticeably across samples. For instance, in the case of the pressure field, the maximum error occurs near the cylinder surface, with approximate values of 0.035, 0.04, 0.05, 0.06, and 0.07 Pa, as visualized in Fig. 15. In Sect. 4.5, we demonstrate that variations in these pressure values ultimately influence the predicted lift and drag forces induced by the pressure field. By sampling within the frameworks of Flow Matching PointNet and Diffusion PointNet, we estimate the maximum potential prediction error and assess model reliability. Consequently, when these models are applied in engineering design, such uncertainty estimates inform the selection of appropriate safety factors.
@@
Similarly, Figure 16 visualizes five samples generated by the Diffusion PointNet model. Differences among these samples are evident, particularly in the maximum absolute errors of both the velocity and pressure fields. For example, examining the generated $u$-velocity reveals maximum errors of approximately $0.04$, $0.20$, $0.25$, and $0.35$ m/s. Notably, in the second row of Fig. 16, or equivalently the sixth row of the same figure, the maximum absolute error of 0.35 occurs at only a single point among all 1024 points in the point cloud, while the remaining points exhibit near-zero absolute error. From a computer graphics perspective, this phenomenon resembles labeling inconsistencies encountered in part‐segmentation tasks when using PointNet [34, 38]. In such cases, more advanced models such as PointNet++ [35], DGCNN [65], and PointMLP [66] have been proposed and could potentially be employed in future studies for generative modeling in computational physics. These models place greater emphasis on capturing local features, thereby reducing the occurrence of outlier-like labels. Hence, these observations provide motivation for exploring more sophisticated point-cloud–based architectures for generative modeling.
@@
Table 3: Comparison between the performance of the Flow Matching PointNet, Diffusion PointNet, and PointNet models for predicting pressure drag and pressure lift on a test set containing 222 unseen geometries. Lift and drag are computed from the predicted pressure fields. Number of samples is equal to 100. In numbers reported in the form $\alpha_{1}\pm\alpha_{2}$, $\alpha_{1}$ and $\alpha_{2}$ show the mean error and the corresponding standard deviation, respectively.
Flow Matching PointNet Diffusion PointNet PointNet
@@
Fig. 17: An example of velocity and pressure distributions predicted by the Flow Matching PointNet, Diffusion PointNet, and PointNet models along the surface of a cylinder with an elliptical cross-section from the test set. The angle is measured counterclockwise from the positive $x$-axis by connecting each surface point to the center of the ellipse at (8, 16).
@@ 4.5 Computation of pressure drag and lift forces
In addition to performing an error analysis over the entire computational domain, we now focus specifically on evaluating the predictions of the velocity and pressure fields along the surface of the object. A comparative analysis is conducted to determine which of the three models, Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, achieves superior performance. In particular, Figure 17 presents the velocity and pressure profiles on the surface of one test object, which corresponds to a cylinder with an elliptical cross-section. As illustrated in Fig. 17, since the flow is viscous, the velocity components in both the $x$- and $y$-directions (i.e., $u$ and $v$) must vanish at the surface. It is evident from Fig. 17 that the predictions obtained by Flow Matching PointNet and Diffusion PointNet are more accurate than those of PointNet for both the velocity and pressure fields. When comparing Flow Matching PointNet and Diffusion PointNet, both models demonstrate satisfactory performance; however, the results in Fig. 17 indicate that Flow Matching PointNet provides slightly more accurate predictions of the zero-velocity condition in the $y$-direction compared to Diffusion PointNet.
@@
Finally, as discussed in Sect. 4.1, 100 samples are drawn by injecting random noise for each geometry from the test set for both Flow Matching PointNet and Diffusion PointNet. Accordingly, the values reported in Table 3 in the form of $\alpha_{1}\pm\alpha_{2}$ indicate $\alpha_{1}$ as the average value and $\alpha_{2}$ as the standard deviation. This representation enables the quantification of uncertainty in the predictions, which is particularly valuable for design-oriented applications. For instance, the average over the maximum absolute error of drag for Flow Matching PointNet is approximately $13.25\%$, accompanied by a standard deviation of $2.14\%$. Similar computations can be performed for Diffusion PointNet and, when relevant, for the maximum absolute error of lift. In contrast, the baseline PointNet model does not provide such uncertainty intervals for error measurements, which represents a key limitation of the model. In summary, both Flow Matching PointNet and Diffusion PointNet not only deliver more accurate predictions but also offer valuable uncertainty information about the model outputs that can assist design engineers, especially when safety margins or worst-case design scenarios must be considered.
Table 4: Robustness test of Flow Matching PointNet, Diffusion PointNet, and PointNet by evaluating the error in predicted velocity and pressure when a certain percentage of points is randomly removed from point clouds in the test set, containing 222 unseen geometries. The $L^{2}$ norm is indicated by $\cdots||$.

### v2

@@ 4.1 Overall results
In this section, we present a comprehensive analysis of the results using tables and histograms. Table 2 reports the relative pointwise errors ($L^{2}$ norm) for predicting the velocity components $u$ and $v$ and the pressure $p$ using the three models: Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. The results are computed across the test set, which contains 222 unseen geometries. Specifically, for each geometry in the test set, 100 samples are generated using both the Flow Matching PointNet and Diffusion PointNet models, meaning that 100 random noise realizations are conditioned on the same geometry. The reported values in Table 2 are expressed in the form $\alpha_{1}\pm\alpha_{2}$, where $\alpha_{1}$ denotes the mean error and $\alpha_{2}$ represents the corresponding standard deviation. According to the values listed in Table 2, the average relative pointwise errors ($L^{2}$ norm) over the 222 geometries for the $u$-velocity component are 1.40%, 1.26%, and 3.20% for the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models, respectively. For the $v$-velocity component, the corresponding errors are approximately 4.81%, 6.17%, and 13.28%. These results demonstrate that the Flow Matching and Diffusion models, particularly for the $v$ variable, achieve substantially lower prediction errors. In comparison with the baseline PointNet, they are capable of reducing the relative pointwise error ($L^{2}$ norm) to below 10%, which is a remarkable improvement from an engineering standpoint. Predicting the $v$ variable is more challenging than predicting $u$ because the flow field exhibits both positive and negative velocities in the $y$-direction due to vortex formation behind the cylinder. Although the $u$ variable may also become negative in some regions, its spatial gradients are typically less steep than those of the $v$ variable.
@@
Examining the results for the pressure variable ($p$) reveals a similar trend. Both the Flow Matching PointNet and Diffusion PointNet models reduce the average relative pointwise error ($L^{2}$ norm) to below 10%, whereas the baseline PointNet exceeds this limit. According to Table 2, the relative pointwise errors ($L^{2}$ norm) for predicting the pressure field are approximately 3.67%, 3.29%, and 10.83% for the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, respectively. The accuracy of the pressure field is particularly important because it directly influences the computation of lift and drag forces, which will be discussed in detail in Sect. 4.5. Furthermore, the uncertainty quantification capabilities of the proposed generative Flow Matching PointNet and Diffusion PointNet models provide valuable insights for practical engineering applications where safety factors must be considered in design processes. Finally, it is worth emphasizing that, as reported in Table 2, the number of trainable parameters, and consequently the model size, is identical across all three frameworks.
@@
In addition to reporting the mean, minimum, and maximum errors across the 222 test geometries in Table 2, the distribution of these errors is illustrated in the histogram shown in Fig. 4. As observed in Fig. 4, Flow Matching PointNet and Diffusion PointNet demonstrate higher accuracy than the baseline PointNet in predicting the $v$ and $p$ variables. However, in all three models, there exist a few outlier cases for which the prediction error increases noticeably, as can be seen in Fig. 4. When comparing the Flow Matching PointNet and Diffusion PointNet in terms of prediction accuracy, the difference between the two is not substantial. Based on the overall results presented in Table 2 and Fig. 4, the Flow Matching PointNet model predicts the $v$-velocity component on average about 1.3% more accurately than the Diffusion PointNet model.
@@
Fig. 5: Evolution of velocity and pressure fields during the reverse process of Flow Matching PointNet at $n=0$ (i.e., $n=N_{\tau}$, standard Gaussian noise), $n=500$, $n=800$, $n=950$, $n=980$, and $n=1000$ (i.e., the clean field).
@@ 4.2 Reverse generative process
Figures 5 and 6 respectively illustrate the reverse processes of the Flow Matching PointNet and the Diffusion PointNet for fluid flow fields around a cylinder with a rectangular cross section, shown as a representative example from the test set. More specifically, Figure 5 presents the velocity and pressure fields at different reverse steps, including $n=0$ (i.e., where all three fields are sampled from standard Gaussian noise), $n=500$, $n=800$, $n=950$, $n=980$, and $n=1000$, which corresponds to the predicted clean velocity and pressure fields. As can be seen from Fig. 5, at $n=500$, the overall structure of the velocity and pressure fields is still largely visually obscured; however, the magnitudes of the variables are approximately within the range of [3, 3][-3,\,3], compared to the case $n=0$, where the variables lie roughly in the range of [6, 6][-6,\,6] due to the standard Gaussian noise. Similarly, at $n=800$, the noise level decreases such that the flow patterns become partially visible. At $n=950$, the velocity and pressure fields are largely reconstructed, and during the remaining reverse steps, they are further refined, with the predicted values converging to physically acceptable ranges. Quantitative error metrics have already been reported in Table 2.
Figure 6 shows similar trends for the reverse process of the Diffusion PointNet, starting from $t=1000$ (i.e., the standard Gaussian noisy field) and progressing to $t=0$, which corresponds to the predicted velocity and pressure fields. By comparing the outputs shown in Figures 5 and 6, it can be observed that the underlying mechanisms of the Flow Matching PointNet and the Diffusion PointNet differ. For example, consider the velocity and pressure field predictions obtained by the Flow Matching PointNet at $n=950$ and by the Diffusion PointNet at $t=50$, corresponding to the same number of reverse steps in the two models. Given the results shown in Fig. 5 and Fig. 6, a visual comparison indicates that the fluid flow patterns obtained by the Flow Matching PointNet exhibit a lower noise level and more closely resemble the predicted clean flow fields at this step (i.e., $n=950$) than those obtained by the Diffusion PointNet at the corresponding step (i.e., $t=50$). This qualitative observation highlights a fundamental difference in the reverse dynamics of the two models.
@@ 4.3 Visual evaluation
Following the numerical analysis, a visual comparison of the results is conducted. Figures 710 present four examples of the Flow Matching PointNet model predictions for velocity and pressure fields, alongside the corresponding ground truth and absolute pointwise error distributions. These figures include cylinders with pentagonal, rectangular, and elliptical cross-sections. Similarly, Figures 1114 illustrate the results of the Diffusion PointNet model for cylinders with hexagonal, triangular, rectangular, and elliptical cross-sections. In all cases, the predictions demonstrate excellent accuracy and a strong agreement with the ground truth. The majority of the absolute errors occur on or near the cylinder surface, where the geometric variations among different test cases are most significant. In Fig. 9 and Fig. 13, the $x$ component of the velocity vector ($u$) becomes negative, indicating the formation of secondary flow regions, and both the Flow Matching PointNet and Diffusion PointNet models successfully capture these phenomena. It should be noted that the test set is identical for all three models: Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. For visual diversity, however, we intentionally select cylinders with different cross-sectional shapes to provide a more comprehensive visual representation in Figs. 714. These results highlight the models’ success in handling diverse geometries with varying shapes, orientations, and sizes. Another notable observation is that the computational domain size differs for each point cloud. For example, as can be seen in Fig. 7, the approximate minimum and maximum $x$-coordinates are 5 m and 11 m, respectively, while the corresponding $y$-coordinates range from 13 m to 19 m. In contrast, as shown in Fig. 9, the $x$-coordinates vary approximately from 2 m to 14 m, and the $y$-coordinates from 10 m to 22 m. A similar comparison can be made between Fig. 11 and Fig. 14. This adaptability reflects a fundamental strength of the PointNet architecture, which can process irregular geometries with varying domain sizes, unlike conventional CNNs that require fixed domain dimensions across the entire dataset. Additionally, the prediction of $\boldsymbol{f}_{\text{target}}$ by Flow Matching PointNet or $\boldsymbol{y}_{\text{noisy}}$ by Diffusion PointNet at each point depends on both the spatial coordinates of that point and the global geometry of the point cloud that contains it. Since the spatial ranges of the $x$ and $y$ coordinates vary across different point clouds in the dataset, accurately predicting $\boldsymbol{f}_{\text{target}}$ or $\boldsymbol{y}_{\text{noisy}}$ is a challenging task. The results presented here demonstrate that the proposed Flow Matching PointNet and Diffusion PointNet models successfully handle this challenge.
@@ 4.4 Sampling diversity
As discussed in Sect. 3, when standard Gaussian noise conditioned on a given geometry (point cloud) is provided as input to the Flow Matching PointNet and Diffusion PointNet models, they generate distinct realizations of the velocity and pressure fields. Here, we compare several of these generated samples. Figure 15 presents five samples generated by the Flow Matching PointNet model. To better illustrate the differences among the generated samples, instead of visualizing the velocity and pressure fields directly, we show the absolute pointwise error along with the corresponding ground truth. As can be seen from Fig. 15, it is evident that each sample differs from the others, even though the regions with the highest errors appear in roughly the same locations. However, the maximum absolute error for the $y$ component of the velocity vector ($v$) and the pressure field ($p$) still varies noticeably across samples. For instance, in the case of the pressure field, the maximum absolute error occurs near the cylinder surface, with approximate values of 0.035, 0.04, 0.05, 0.06, and 0.07 Pa, as visualized in Fig. 15. In Sect. 4.5, we demonstrate that variations in these pressure values ultimately influence the predicted lift and drag forces induced by the pressure field. By sampling within the frameworks of Flow Matching PointNet and Diffusion PointNet, we estimate the maximum potential prediction error and assess model reliability. Consequently, when these models are applied in engineering design, such uncertainty estimates inform the selection of appropriate safety factors.
@@
Similarly, Figure 16 visualizes five samples generated by the Diffusion PointNet model. Differences among these samples are evident, particularly in the maximum absolute errors of both the velocity and pressure fields. For example, examining the generated $u$-velocity reveals maximum errors of approximately $0.04$, $0.20$, $0.25$, and $0.35$ m/s. Notably, in the second row of Fig. 16, or equivalently the sixth row of the same figure, the maximum absolute error of 0.35 occurs at only a single point among all 1024 points in the point cloud, while the remaining points exhibit near-zero absolute error. From a computer graphics perspective, this phenomenon resembles labeling inconsistencies encountered in part‐segmentation tasks when using PointNet [34, 38]. In such cases, more advanced models such as PointNet++ [35], DGCNN [66], and PointMLP [67] have been proposed and could potentially be employed in future studies for generative modeling in computational physics. These models place greater emphasis on capturing local features, thereby reducing the occurrence of outlier-like labels. Hence, these observations provide motivation for exploring more sophisticated point cloud-based architectures for generative modeling.
@@
Table 3: Comparison between the performance of the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models for predicting pressure drag and pressure lift on a test set containing 222 unseen geometries. Lift and drag are computed from the predicted pressure fields. Number of samples is equal to 100. In numbers reported in the form $\alpha_{1}\pm\alpha_{2}$, $\alpha_{1}$ and $\alpha_{2}$ show the mean error and the corresponding standard deviation, respectively.
Flow Matching PointNet Diffusion PointNet Baseline PointNet
@@
Fig. 17: An example of velocity and pressure distributions predicted by the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models along the surface of a cylinder with an elliptical cross-section from the test set. The angle is measured counterclockwise from the positive $x$-axis by connecting each surface point to the center of the ellipse at (8, 16).
@@ 4.5 Computation of pressure drag and lift forces
In addition to performing an error analysis over the entire computational domain, we now focus specifically on evaluating the predictions of the velocity and pressure fields along the surface of the object. A comparative analysis is conducted to determine which of the three models, Flow Matching PointNet, Diffusion PointNet, and baseline PointNet, achieves superior performance. In particular, Figure 17 presents the velocity and pressure profiles on the surface of one test object, which corresponds to a cylinder with an elliptical cross-section. As illustrated in Fig. 17, since the flow is viscous, the components of the velocity vector (i.e., $u$ and $v$) must vanish at the surface. It is evident from Fig. 17 that the predictions obtained by Flow Matching PointNet and Diffusion PointNet are more accurate than those of baseline PointNet for both the velocity and pressure fields. When comparing Flow Matching PointNet and Diffusion PointNet, both models demonstrate excellent performance; however, the results shown in Fig. 17 indicate that Flow Matching PointNet provides slightly more accurate predictions of the zero-velocity condition in the $y$-direction compared to Diffusion PointNet.
@@
Finally, as discussed in Sect. 4.1, 100 samples are drawn by injecting standard Gaussian noise for each geometry from the test set for both Flow Matching PointNet and Diffusion PointNet. Accordingly, the values reported in Table 3 in the form of $\alpha_{1}\pm\alpha_{2}$ indicate $\alpha_{1}$ as the average value and $\alpha_{2}$ as the standard deviation. This representation enables the quantification of uncertainty in the predictions, which is particularly valuable for design-oriented applications. For instance, the average over the maximum absolute error of drag for Flow Matching PointNet is approximately $13.25\%$, accompanied by a standard deviation of $2.14\%$, as listed in Table 3. Similar information for the maximum absolute error of lift using Diffusion PointNet is provided in Table 3. In contrast, the baseline PointNet model does not provide such uncertainty intervals for error measurements, which represents a key limitation of the model. In summary, both Flow Matching PointNet and Diffusion PointNet not only deliver more accurate predictions but also offer valuable uncertainty information about the model outputs that can assist design engineers, especially when safety margins or worst-case design scenarios must be considered.
Table 4: Robustness test of Flow Matching PointNet, Diffusion PointNet, and baseline PointNet by evaluating the error in predicted velocity and pressure fields when a certain percentage of points is randomly removed from point clouds in the test set, containing 222 unseen geometries. The $L^{2}$ norm is indicated by $\cdots||$.